# Open Domain Question Answering with Conflicting Contexts

### v1

@@ Abstract
Open domain question answering systems frequently rely on information retrieved from large collections of text (such as the Web) to answer questions. However, such collections of text often contain conflicting information, and indiscriminately depending on this information may result in untruthful and inaccurate answers. To understand the gravity of this problem, we collect a human-annotated dataset, Question Answering with Conflicting Contexts (QACC), and find that as much as $25\%$ of unambiguous, open domain questions can lead to conflicting contexts when retrieved using Google Search. We evaluate and benchmark three powerful Large Language Models (LLMs) with our dataset QACC and demonstrate their limitations in effectively addressing questions with conflicting information. To explore how humans reason through conflicting contexts, we request our annotators to provide explanations for their selections of correct answers. We demonstrate that by finetuning LLMs to explain their answers, we can introduce richer information into their training that guide them through the process of reasoning with conflicting contexts. We will release our dataset and code to promote research along this line.
@@
Siyi Liu thanks: Work done during internship at AWS AI Lab 1,2, Qiang Ning1, Kishaloy Halder1, Wei Xiao1, Zheng Qi1, Phu Mon Htut1, Yi Zhang1, Neha Anna John1, Bonan Min1, Yassine Benajiba1, Dan Roth1,2, 1AWS AI Lab, 2University of Pennsylvania, siyiliu@seas.upenn.edu
@@ 1 Introduction
To quantify how often conflicting contexts occur on the web, we construct our dataset named QACC (Question Answering with Conflicting Contexts). We consider unambiguous open domain questions from AmbigQA Min et al. (2020) and use Google Search API [^2] https://developers.google.com/custom-search/v1/overview to retrieve up to $10$ results for each question. We then use Amazon Mechanical Turk and ask human annotators to determine whether there exists different answers in the contexts. We find that about $25\%$ of the unambiguous open domain questions will yield conflicting evidence from Google. We evaluate three popular LLMs (GPT-3.5, Claude-3, and Phi-3) on our dataset with different prompting and finetuning strategies and establish that conflicting contexts can lead to substantial performance degradation in them. To understand how humans reason through conflicting contexts, we ask our annotators to select from a pre-defined set of reasons when deciding on the answer. Our findings indicate that humans often adhere to majority vote (i.e. selecting the most popular answer) when seeing conflicting contexts. In addition, we also request our annotators to provide a single sentence, natural language explanation for their answers. We find that by finetuning LLMs to explain their answers, we can introduce richer information into their training that guide them through the process of reasoning with conflicting contexts and improve their performance in both QACC and a perturbed NQ-Open dataset Lee et al. (2019).
@@ 3.1 Problem Definition
Given a question $q$, a list of retrieved contexts $C=\{c_{1},c_{2},...,c_{i}\}$, and a question answering system $\phi$, we can get a list of individual answers $A=\{a_{1},a_{2},...,a_{i}\}$, where $a_{i}=\phi(q,c_{i})$. We state that the question $q$ has conflicting evidence if and only if $\exists(a_{i},a_{j}\in A)(a_{i}\neq a_{j})$. In other words, at each step a question answering system (a human or a language model) is given the question and only one of the context in order to answer the question. Iterate this through all of the contexts, and we state that there are conflicting contexts if and only if there are different answers generated when given different contexts. We believe that this definition over the space of the answers makes it clearer and more well-defined to instruct the human annotators to determine if there are conflicting contexts.
@@ 3.3 Human Annotation
We employ a rigorous human annotation pipeline with a qualification exam before the main annotation task, a strategy commonly used to ensure the collection of high-quality datasets Han et al. (2021); Dasigi et al. (2021). Only annotators that have passed our qualification exams can participate in the main annotation task. We use Amazon Mechanical Turk to collect the annotations and CROWDAQ to design the annotation interfaceNing et al. (2020). Each question is annotated by one annotator and paid with \$$\$0.35$ in the main annotation task. Examples of our qualification and annotation interfaces are shown in Appendix A.
@@ 3.5 Analysis
We also conduct data analysis to study the types of questions that lead to conflicting contexts more often. In figure 4, we see that the two most prominent types of questions are Why and How questions. 40% of questions in QACC that start with Why and 38.75% of questions that start with How are determined as having conflicting contexts on Google. This aligns with our hypothesis since questions that start with Why and How are typically open-ended questions with a more complex answer and can involve different perspectives.
Figure 4: This shows the types of questions that are labeled by our annotators as having conflicts. Each bar represents the percentage of all of the questions of this type that have conflicting contexts.
@@ 4 Experiments
GPT-3.5-fewshot Context 42.03 53.47 50.55 59.21 68.14 65.87
+ Majority Vote 42.99 56.60 53.13 59.80 69.73 67.20
+ Dis & Ans 42.99 55.44 52.27 58.94 69.01 66.45
+ Exp & Ans 42.03 54.95 51.66 58.13 68.33 65.73
@@ 4.3 Experiment Setup
We conduct experiments on three popular families of instruction-tuned large language models: GPT3.5-turbo-0125, Claude3-Sonnet, and Phi-3-Medium-Instruct (14B), with zero-shot inference, few-shot inference, and finetuning. We find that LLMs greatly benefit from in-context examples (few-shot) compared to zeroshot (See Appendix B) when answering open domain questions, so we only present few-shot inference and finetuning results in Table 2.
For few-shot inference experiments, we include one in-context example of expected input-output pair when prompting the three language models. For finetuning experiments, we finetune Phi-3-Medium-Instruct using LoRA Hu et al. (2021) with language modeling loss (SFT). We first find the best hyperparameters of finetuning using the validation set of QACC and then train on both the training and validation set and report results in the test set. We also use the validation set of QACC to find the best prompt and instruction format for each methods and use them for both fewshot inference and finetuning. We use Exact Match and F1 scores as the metrics for all our evaluations.
4.4 Experiment Results
QACC
Table 2 exhibits our experiment results in QACC. We can see that all LLMs that we evaluate inevitably experience worse performance when there are conflicting contexts, comparing their results on EM-C and EM-NC, as well as their results on F1-C and F1-NC. We also find that in different LLMs, their best prompting methods in the fewshot setting are also different. GPT-3.5 has the best performance when prompted to take majority vote when seeing conflicting contexts, Claude-3 gives the best results when instructed to first explain and then answer the question. Phi-3 presents comparable performances when instructed to take majority vote and explain then answer.
Perturbed NQ-Open
@@ 5 Conclusion
Eliciting natural language explanations from LLMs have several limitations. Previous work have shown that explanations generated by LLMs can be unreliable and can lead to wrong interpretations of the models. However, in this work, we focus on the improvement of reasoning abilities of LLMs when finetuning explanations, rather than making interpretations of their explanations.
Appendix A Human Annotation Interface
Figure 5: Qualification example 1.
Figure 6: Qualification example 1.
Figure 7: Annotation example 1.
Figure 8: Annotation example 2.
Figure 9: Annotation example 3.

### v3

@@ Abstract
Open domain question answering systems frequently rely on information retrieved from large collections of text (such as the Web) to answer questions. However, such collections of text often contain conflicting information, and indiscriminately depending on this information may result in untruthful and inaccurate answers. To understand the gravity of this problem, we collect a human-annotated dataset, Question Answering with Conflicting Contexts (QACC), and find that as much as $25\%$ of unambiguous, open domain questions can lead to conflicting contexts when retrieved using Google Search. We evaluate and benchmark three powerful Large Language Models (LLMs) with our dataset QACC and demonstrate their limitations in effectively addressing questions with conflicting information. To explore how humans reason through conflicting contexts, we request our annotators to provide explanations for their selections of correct answers. We demonstrate that by finetuning LLMs to explain their answers, we can introduce richer information into their training that guide them through the process of reasoning with conflicting contexts. We will release our dataset and code to promote research along this line. Please direct requests and inquires about the data and code to the authors.
@@
Siyi Liu thanks: Work done during internship at AWS AI Labs 1,2, Qiang Ning1, Kishaloy Halder1, Wei Xiao1, Zheng Qi1, Phu Mon Htut1, Yi Zhang1, Neha Anna John1, Bonan Min1, Yassine Benajiba1, Dan Roth1,2, 1AWS AI Labs, 2University of Pennsylvania, siyiliu@seas.upenn.edu
@@ 1 Introduction
To quantify how often conflicting contexts occur on the web, we construct our dataset named QACC (Question Answering with Conflicting Contexts). We consider unambiguous open domain questions from AmbigQA Min et al. (2020) and use Google Search API [^2] https://developers.google.com/custom-search/v1/overview to retrieve up to $10$ results for each question. We then use Amazon Mechanical Turk and ask human annotators to determine whether there exists different answers in the contexts. We find that about $25\%$ of the unambiguous open domain questions will yield conflicting evidence from Google. We evaluate three popular LLMs (GPT-4o, Claude-3, and Phi-3) on our dataset with different prompting and finetuning strategies and establish that conflicting contexts can lead to substantial performance degradation in them. To understand how humans reason through conflicting contexts, we ask our annotators to select from a pre-defined set of reasons when deciding on the answer. Our findings indicate that humans often adhere to majority vote (i.e. selecting the most popular answer) when seeing conflicting contexts. In addition, we also request our annotators to provide a single sentence, natural language explanation for their answers. We find that by finetuning LLMs to explain their answers, we can introduce richer information into their training that guide them through the process of reasoning with conflicting contexts and improve their performance in both QACC and a perturbed NQ-Open dataset Lee et al. (2019).
@@ 3.1 Problem Definition
Given a question $q$, a list of retrieved contexts $C=\{c_{1},c_{2},...,c_{i}\}$, and a question answering system $\phi$, we can get a list of individual answers $A=\{a_{1},a_{2},...,a_{i}\}$, where $a_{i}=\phi(q,c_{i})$. We state that the question $q$ has conflicting evidence if and only if $\exists(a_{i},a_{j}\in A)(a_{i}\neq a_{j})$. In other words, at each step a question answering system (a human or a language model) is given the question and only one of the context in order to answer the question. Iterate this through all of the contexts, and we state that there are conflicting contexts if and only if there are different answers generated when given different contexts. Note that our definition of conflict here differs from the definition of contradiction in traditional Natural Language Inference (NLI) tasks. Here we define conflict in a way that is less restricting than contradictory texts in NLI to further exploit its applicability in our target domain, open domain QA with RAG. For instance, in a RAG scenario, retrievers can often retrieve contexts that contain seemingly correct answers (e.g. neutral texts in the case of NLI), and such different/conflicting answers may also confuse the downstream LLMs. We believe that our definition of conflict and the resulted dataset can therefore better support us towards our goal.
@@ 3.3 Human Annotation
We employ a rigorous human annotation pipeline with a qualification exam before the main annotation task, a strategy commonly used to ensure the collection of high-quality datasets Han et al. (2021); Dasigi et al. (2021). Only annotators that have passed our qualification exams can participate in the main annotation task. We use Amazon Mechanical Turk to collect the annotations and CROWDAQ to design the annotation interfaceNing et al. (2020). Each question is annotated by one annotator and paid with \$$\$0.35$ in the main annotation task. Examples of our qualification and annotation interfaces are shown in Appendix D.
@@ 3.5 Analysis
We also conduct data analysis to study the different types of questions in our dataset that lead to conflicting contexts. In figure 4, we see that most of the questions in our dataset are Who questions. This type of question has about 20% of times that lead to conflicting contexts. How questions, on the other hand, can lead to conflicting contexts almost 40% of the times. This aligns with our hypothesis since questions that start with How are typically open-ended questions with a more complex answer and can involve different perspectives.
Another characteristic about QACC that we observe is that the most prominent type of conflicts between different answers is the mismatch of entities, i.e., different names, time, and places. This is largely due to the fact that we source our questions from AmbigQA and Natural Questions, where the expected answers are short phrases that are mostly entities.
Figure 4: Different types of questions in our dataset that have conflicts.
@@ 4 Experiments
GPT-4o-fewshot Context 42.51 54.13 51.17 59.29 70.06 67.32
+ Majority & Ans 40.58 53.96 50.55 57.35 69.03 66.05
+ Discern & Ans 39.13 52.15 48.83 56.33 67.81 64.88
+ Exp & Ans 38.65 54.29 50.31 57.43 69.19 66.2
@@ 4.3 Experiment Setup
We conduct experiments on three popular instruction-tuned large language models from different families: GPT-4o-mini, Claude3-Sonnet, and Phi-3-Medium-Instruct (14B), with zero-shot inference, few-shot inference, and finetuning. We find that LLMs greatly benefit from in-context examples (few-shot) compared to zeroshot (see Appendix B) when answering open domain questions, so we only present few-shot inference and finetuning results in Table 2.
For few-shot inference experiments, we include one in-context example of expected input-output pair when prompting the three language models. For finetuning experiments, we finetune Phi-3-Medium-Instruct using LoRA Hu et al. (2021) with language modeling loss (SFT). We first find the best hyperparameters of finetuning using the validation set of QACC and then train on both the training and validation set and report results in the test set. We also use the validation set of QACC to find the best prompt and instruction format for each methods and use them for both fewshot inference and finetuning. More details of experiment settings are discussed in Appendix A
We follow the conventions and use Exact Match and F1 scores as the metrics for all our evaluations. Exact Match returns positive if the generated answer is identical to the reference and negative if otherwise, whereas F1 score is more forgiving and measures the word overlap between the generated and reference answers. We note that LLMs are prone to long generations, so we specifically instruct all of the models to answer with as few words as possible in the prompts (see examples of the prompts in Appendix C).
@@
4.4 Experiment Results
QACC
Table 2 exhibits our experiment results in QACC. We can see that all LLMs that we evaluate inevitably experience worse performance when there are conflicting contexts, comparing their results on EM-C and EM-NC, as well as their results on F1-C and F1-NC. We also find that in different LLMs, their best prompting methods in the fewshot setting are also different. GPT-4o has the best performance when prompted with just the contexts when seeing conflicting contexts, Claude-3 gives the best results when instructed to first explain and then answer the question, and Phi-3 presents comparable performances when instructed to take majority vote and explain then answer.
Perturbed NQ-Open
4.5 Analysis
Figure 5 shows the performance of fewshot GPT-4o on the set of questions in QACC that has conflicting contexts. We can see that LLMs like GPT-4o can answer open domain questions reasonably well even with conflicting contexts when the questions are asking about a person (Who types of question). However, GPT-4o fails significantly more when the question is asking about a time (When), about a place (Where), or when the question is open-ended (How). We hypothesize that this may relate to the pretraining corpus of LLMs and the frequency that different entities appear in the pretraining corpus: popular people names that appear frequently in LLMs’ pretraining corpus allow them to utilize their parametric knowledge to distinguish the answers among the conflicting contexts, whereas the different timestamps and places exist more sparsely in the corpus (as well as on the web), making it more difficult for the LLMs to discern.
Figure 5: Fewshot GPT-4o performance on the test set of QACC that has conflicting contexts. The x-axis indicates the different types of questions and the y-axis denotes the F1 score for each type.
@@ 5 Conclusion
Eliciting natural language explanations from LLMs have several limitations. Previous work have shown that explanations generated by LLMs can be unreliable and can lead to wrong interpretations of the models. However, in this work, we focus on the improvement of reasoning abilities of LLMs when finetuning with explanations data, rather than making interpretations of their explanations.
Appendix A Experiment Details
We perform our finetuning experiments using Amazon Elastic Compute Cloud (Amazon EC2). We use one p4d.24xlarge instance for the training. It has 8 NVIDIA A100 GPUs with 40.0 GB GPU memory each. We use Parameter-Efficient Fine-Tuning (PEFT) and Low-Rank Adaptation (LoRA) to train our models. We use learning rata 2e-4, lora_r = 16, lora_alpha = 32, and lora_dropout = 0.05.
We use snippets returned by Google Search as the contexts for each of the articles retrieved. Studies have shown that LLMs struggle to process and understand very long contexts Liu et al. (2023), so we believe that Google snippets are good summarization and extraction of their original articles for our model to process different contexts. To guarantee that the answer always exists in the contexts, we remove examples that our annotators believe they didn’t find a correct answer in the contexts in our experiments. Examples of our annotation interface are shown in Appendix D.
@@ Appendix A Experiment Details
Appendix C Example Prompt
We present an example of the prompt and in-context example we use for running our experiments. We preserve the same format and instruction in our finetuning experiments for consistency.
Instruction
Given the following contexts, explain your reasoning first and then provide a direct, concise answer to the question.
Make the answer as short as possible and use \ n \ n as the delimiter to distinguish your explanation and answer.
In-context example
Context1 from www.sho.com: Kerris Lilla Dorsey is best known for her roles as Brad Pitt’s daughter in the
Oscar® nominated film MONEYBALL and as Steve Carell and Jennifer Garner’s ... Context2 from www.imdb.com: She is known for her roles as Paige Whedon in the television series Brothers &
Sisters, Casey Beane, Billy Beane’s (Brad Pitt) daughter, in the 2011 film ... Context3 from people.com: Jul 2, 2015 ... Kerris Dorsey has worked with several of Hollywood’s most
famous dads. She currently plays Liev Schreiber’s daughter on Ray Donovan and got ... Context4 from www.imdb.com: She is known for her roles as Paige Whedon in the television series Brothers & Sisters,
Casey Beane, Billy Beane’s (Brad Pitt) daughter, in the 2011 film ... Context5 from www.sportskeeda.com: Feb 11, 2023 ... All about Casey Beane, who was pictured in Moneyball as a young
girl ... The 2011 film Moneyball was one of the most popular baseball movies ever ... Context6 from en.wikipedia.org: Moneyball is a 2011 American biographical sports drama film directed by Bennett
Miller with ... In the film, Beane (Brad Pitt) and assistant general manager Peter Brand ... Context7 from www.latimes.com: Sep 10, 2011 ... And “Moneyball” is all about Brad Pitt. ... And while you get
a couple of glimpses of Pitt’s daughter, played by 13-year-old Kerris Dorsey, ... Context8 from www.pinterest.com: Jan 9, 2020 ... ... (played by Brad Pitt) daughter, in the 2011 film Moneyball,
... Dorsey plays Bridget Donovan, the daughter of the title character, ...
Who plays brad pitt’s daughter in moneyball?
Explanation: I chose it because it is the only correct answer and it is explicitly mentioned in context 4 and 5.
Answer: Kerris Lilla Dorsey
Appendix D Human Annotation Interface
Figure 6: Qualification example 1.
Figure 7: Qualification example 1.
Figure 8: Annotation example 1.
Figure 9: Annotation example 2.
Figure 10: Annotation example 3.